OpenStudy (anonymous):

i simplified (6x^9/4x^4) x (12x^2/3x^5) to 6/1 but i dont know what to do with the exponents. please help

OpenStudy (anonymous):

\[(6x ^{9}/4x ^{4}) \times (12x ^{2}/3x ^{5})\]

OpenStudy (anonymous):

6x^9/4x^4 = 6x^(5)/4 = 3x^(5)/2 12x^(2)/3x^(5) = 12/3x^(3) = 4/x^(3) so (3x^(5)/2) * (4/x^(3)) = 4(3x^(5))/x^(3)(2) = 12x^(5)/2x^(3) = 6x^(2) Here are the steps remember that x^(-1) = 1/x so 6x^9/4x^4 = 6x^(9)x^(-4) 9-4 = 5 therefore 6x^(5)/4

OpenStudy (anonymous):

the x^(-1) = 1/x is a very important rule that you will need to know for higher math :)

OpenStudy (anonymous):

I made a mistake 6x^9/4x^4 = 6x^(9)x^(-4)/4

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Remember when multiplying exponents it is additive. for example (1 + x)^(2) = (1+x)(1+x) its just an easier way to write it but if we went (1 + x)^(2) * (1 + x)^(2) we would just have (1+x)(1+x)(1+x)(1+x) so it is an addditive process. Now if we consider the rule that x^(-1) = 1/x (1+x)/(1+x) = (1+x)^(1) * (1+x)^(-1) = (1+x)^(0) = 1 as anything to the power of 0 equals one excluding 0^(0) which is indeterminate

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

I hope that you read what I wrote because knowing this will really strengthen your skills with algebra

OpenStudy (anonymous):

no i dont understand. all you did was confuse me

OpenStudy (anonymous):

so understand that (x) = (x)^(1) they are equivilent same with (x+1) = (x+1)^(1) you can raise it to the power of one. When mutilplying two like variables exponents are another way of showing how many times you are multiplying the same variable. So if we have x x^(1) = x but if we had x^(2) = x * x (x * x is the same as saying (x^(1) * x^(1)) so 1 + 1 = 2 thus x^(2) do you understand so far?

OpenStudy (anonymous):

Please stay with me as I'm doing this as a study break and I need to get back to work very soon, if you want me to explain this important concept to you

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

x*x*x = x^(3) same as saying x^(1) * x^(1) * x^(1) = x^(3) 1 + 1 + 1 = 3

OpenStudy (anonymous):

yeah i know that concept. but what do i do with my exponents to simplify them?

OpenStudy (anonymous):

Ok there is a rule x^(-1) = 1/x

OpenStudy (anonymous):

remember this rule

OpenStudy (anonymous):

so taking in account that rule x^(1)/x^(1) = x^(1) * x^(-1) = x^(0) = 1 1 + (-1) = 0 remembering that + * - = - Also remember that anything to the power of 0 is equal to 1 (excluding 0 which you can't raise to the power of 0) 5^(0) = 1 y^(0) = 1 etc

OpenStudy (anonymous):

so to 6x^(6)/4x^(4) = (6x^(6) * 4x^(-4))/4 = 6x^(2)/4 6 + (-4) = 6 - 4 = 2

OpenStudy (anonymous):

12x^(2)/3x^(5) = 12x^(2)x^(-5)/3 = 12x^(-3)/3 = (12/3)(1/x^(3)) = 12(1)/3x^(3) = 12/3x^(3) 2 + (-5) = 2 - 5 = -3 x^(-3) = 1/x^(3)

OpenStudy (anonymous):

I hope this helps and you are just not even more confused it is some what logical if you jsut accept the rule that x^(-y) = 1/x^(y) y being any number

OpenStudy (anonymous):

If you are still lost tell me what you are not getting and I will try to explain it better. Which will be kind of hard because there is only so many ways to explain a concept

OpenStudy (anonymous):

Also remember that this can be only done with like variables or constants (like 2) 2/2 = 2^(1)/2^(1) = 2^(1)2^(-1) = 2^(0) = 1 1 + (-1) = 1 - 1 = 0

OpenStudy (anonymous):

also remember that larger numbers can represent smaller numbers 8/2 = 2/2*2 = 2^(3) = 8 = 2^(3)/2^(1) = 2^(3) * 2^(-1) = 2^(2) = 4 3 + (-1) = 3 - 1 = 2

OpenStudy (anonymous):

this sort of explains division I find

OpenStudy (anonymous):

sorry 2*2*2 = 2^(3) = 8 but yeah I will wait for you to respond.

OpenStudy (anonymous):

Or you can just give up and take the answer, and I can just accept I wasted my time lol

OpenStudy (anonymous):

no you didnt lol. i figured it out

OpenStudy (anonymous):

Great :)